Calculate the acceleration of a car if it is moving from 20m/s to 50m/s in 6sec.

a. 5 m/s2 b. 6 m/s2 c. 11.67 m/s2 d. 15 m/s2




1
Expert's answer
2021-04-19T17:09:22-0400

The acceleration of the car is "a = \frac{v_1 - v_0}{t} = \frac{50 \frac{m}{s} - 20 \frac{m}{s}}{6 s} = 5 \frac{m}{s^2}", hence the answer is a).